A jeweler has five rings, each weighing 18grams, made of an alloy of 10% silver and 90% gold. He decides to melt down the rings and add enough silver to reduce the gold content to 60%. How much silver should he add?
Answer by ankor@dixie-net.com(22740)

A jeweler has five rings, each weighing 18 grams, made of an alloy of 10% silver and 90% gold.
He decides to melt down the rings and add enough silver to reduce the gold content to 60%.
How much silver should he add?
:
60% gold means it will have 40% silver
total weight of the 5 rings: 5*18 = 90 grams
let s = amt of silver required
write an equation base on the percent silver
.10(90) + s = .40(90+s)
9 + s = 36 + .40s
s - .40s = 36 - 9
.6s = 27
s = 27/.6
s = 45 grams of silver required